Hi ... I was hoping someone can help me with this problem.
The population of Springfield is decreasing at a rate of 250 people per year. Find the data that is missing from the chart.
Year Population
? 8400
2002 8150
2008 ?
? 5900
? 4900
Thanks so much!!

Answer by edjones(8007)

Year Population
2001 8400
2002 8150
2008 6650
2011 5900
2015 4900
.
2008-2002=6
6*-250=-1500
8150-1500=6650
.
6650-5900=750
750/250=32008+3=2011
.
5900-4900=1000
1000/250=4
2011+4=2015
.
8400-8150=250
250/250=1
2002-1=2001
.
Ed